OpenStudy (kearsten):

x^2=10x FAN AND MEDAL

OpenStudy (anonymous):

x^2=10x (x^2=10x)/x x=10

OpenStudy (cookie_2046):

Yep =10

OpenStudy (anonymous):

;)